OpenStudy (anonymous):

if first term of a G.P. is 5, the sum of the first 3 terms is 31/5 find the common ratio.

Parth (parthkohli):

\[5 + 5r + 5r^2 = \frac {31}{5}\]A quadratic equation. Find the two solutions and see which one works.

OpenStudy (anonymous):

yes got it thanks :)

OpenStudy (amistre64):

i spose if you want a simple solution that would be fine ... :)

OpenStudy (amistre64):

i was thinking of: \[5\frac{1-r^3}{1-r}=\frac{31}{5}\]

Parth (parthkohli):

honestly i didn't use the sum formula because i couldn't remember it myself :P

OpenStudy (amistre64):

1-r^3 = (1-r)(1+r+r^2) so it reduces that same